56 litres of oxygen gas combines with 112litres of hydrogen to form the water when 336 litres hydrogen passed over heated cupric oxide ,the latter loses 0.24grms of its weight. if all mesurements of volume are done at s.t.p prove that data illustrates the law of definate proprations.

Grade:12th Pass

1 Answers

Aman Bansal
12 years ago

Dear Ravi,

The Law of Multiple Proportions is the third postulate of Daltons atomic theory. It states that the masses of one element which combine with a fixed mass of the second element are in a ratio of whole numbers.

Therefore, the masses of oxygen in the two compounds that combine with a fixed mass of carbon should be in a whole-number ratio. In 100 g of the first compound (100 is chosen to make calculations easier) there are 57.1 g O and 42.9 g C. The mass of O per gram C is:

57.1 g O / 42.9 g C = 1.33 g O per g C

In the 100 g of the second compound, there are 72.7 g O and 27.3 g C. The mass of oxygen per gram of carbon is:

72.7 g O / 27.3 g C = 2.66 g O per g C

Dividing the mass O per g C of the second (larger value) compound:

2.66 / 1.33 = 2

Which mean that the masses of oxygen that combine with carbon are in a 2:1 ratio. The whole-number ratio is consistent with the Law of Multiple Proportions.

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert